find the exact solution for x^2+6x+29=0

Answer by TimothyLamb(4379)

xx + 6x + 29 = 0
---
the above quadratic equation is in standard form, with a=1, b=6 and c=29
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 6 29
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two complex roots at:
---
x = -3 + (4.47213595)i
x = -3 - (4.47213595)i
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---